# Eigenmatrix for unstructured sparse recovery

### 09 Feb 2024 v2

@@ Abstract.
This paper considers the unstructured sparse recovery problems in a general form. Examples include rational approximation, spectral function estimation, Fourier inversion, Laplace inversion, and sparse deconvolution. The main challenges are the noise in the sample values and the unstructured nature of the sample locations. This paper proposes the eigenmatrix, a data-driven construction with desired approximate eigenvalues and eigenvectors. The eigenmatrix offers a new way for these sparse recovery problems. Numerical results are provided to demonstrate the efficiency of the proposed method.
@@
Sparse recovery, Prony’s method
@@
This work is partially supported by NSF grant DMS-2011699 and DMS-2208163.
@@ 1. Introduction
This paper considers the unstructured sparse recovery problems of a general form. Let $X$ be the parameter space, typically a subset of $\mathbb{R}$ or $\mathbb{C}$, and $S$ be the sampling space. $G(s,x)$ is the kernel function for $s\in S$ and $x\in X$, and is assumed to be analytic in $x$. Suppose that
@@
  • Laplace inversion. $G(s,x)=\exp(-sx)$, $X$ is an interval $[c_{1},c_{2}]$ of the positive real axis, and $\{s_{j}\}$ is a set of positive real numbers.
@@
The primary challenges of the current setup come from two sources. First, the sample values $\{\tilde{u}_{j}\}$ are noisy, which makes the recovery problem quite ill-posed when the kernel $G(s,x)$ is numerically low-rank. Second, the sample locations $\{s_{j}\}$ are unstructured, which excludes many existing algorithms that exploit special structures.
@@ 1.1. Contribution
This paper introduces the eigenmatrix for these unstructured sparse recovery problems. By defining the vector-valued function $\mathbf{g}(x)=[G(s_{j},x)]_{1\leq j\leq n_{s}}$ for $x\in X$, we introduce the eigenmatrix as an $n_{s}\times n_{s}$ matrix $M$ that satisfies, for any $x\in X$
@@ 1.2. Related work
The rest of the paper is organized as follows. Section 2 describes the eigenmatrix construction. Section 3 presents the numerical experiments of the applications mentioned above. Section 4 concludes with a discussion for future work.
2. Algorithm
@@
To motivate the eigenmatrix construction, we first briefly review Prony’s method. Consider the line-spectrum estimation (or superresolution) problem with $G(s,x)=\exp(2\pi isx)$, $X=[0,1]$, and $s_{j}=j$ for $j\in\mathbb{Z}$. Here, we make the simplifying assumption that $\{s_{j}\}$ is the whole integer lattice, though only a finite chunk is required in the actual Prony’s method.
@@
Introduce the infinitely-long vector-valued function $\mathbf{g}(x)=[G(s_{j},x)]_{j\in\mathbb{Z}}=[\exp(2\pi ijx)]_{j\in\mathbb{Z}}$. Let $M$ be the shifting matrix that moves each entry up by one slot, i.e., $(M\mathbf{v})_{k}=\mathbf{v}_{k+1}$ for any vector $\mathbf{v}$. Then
$M\mathbf{g}(x)=e^{2\pi ix}\mathbf{g}(x)\quad\text{or equivalently}\quad(M-e^{2\pi ix}I)\mathbf{g}(x)=0.$
Since $(M-e^{2\pi ix_{\ell}}I)\mathbf{g}(x_{\ell})=0$ for each $\ell$ and the operators $M-e^{2\pi ix_{\ell}}I$ for different $x_{\ell}$ commute with each other, we have
$\prod_{\ell=1}^{n_{x}}(M-e^{2\pi ix_{\ell}}I)\mathbf{g}(x_{k})=0.$
@@
Define the vector $\mathbf{u}=[u_{j}]_{j\in\mathbb{Z}}$ of the exact observations $u_{j}=u(s_{j})=\sum_{k}G(s_{j},x_{k})w_{k}=\sum_{k}e^{2\pi ijx_{k}}w_{k}=\sum_{k}w_{k}\mathbf{g}(x_{k})$. Since $\mathbf{u}$ is a linear combination of $\{\mathbf{g}(x_{k})\}_{1\leq k\leq n_{x}}$, we have
(1) $\prod_{\ell=1}^{n_{x}}(M-e^{2\pi ix_{\ell}}I)\mathbf{u}=0.$
Because $\prod_{\ell}(M-e^{2\pi ix_{\ell}}I)$ is a matrix polynomial of form $p(M):=p_{0}+p_{1}M+\cdots+p_{n_{x}}M^{n_{x}}$, (1) translates to
$(p_{0}+p_{1}M+\cdots+p_{n_{x}}M^{n_{x}})\mathbf{u}=0.$
This is equivalent to
@@
(2) $\begin{bmatrix}\mathbf{u}&M\mathbf{u}&\ldots&M^{n_{x}}\mathbf{u}\end{bmatrix}\mathbf{p}=0\quad\text{with}\quad\mathbf{p}=\begin{bmatrix}p_{0}\\ \vdots\\ p_{n_{x}}\end{bmatrix}$
@@
One can solve for the vector $\mathbf{p}$ by identifying the null space vector of the matrix in (2). The roots of the polynomial $p(x)=p_{0}+p_{1}x+\ldots p_{n_{x}}x^{n_{x}}$ are $\{e^{2\pi ix_{k}}\}$. Given the sample locations $\{x_{k}\}$, the sample weights $\{w_{k}\}$ can be computed, for example via the least square solve.
@@
For most problems, the sample values $\{\tilde{u}_{j}\}$ have noise. As a result, the sample locations and weights obtained above are only approximations. Most implementations of Prony’s method have a postprocessing step, where these approximations are used as the initial guesses of the following optimization problem
(3) $\min_{\tilde{x}_{k},\tilde{w}_{k}}\sum_{j}\left|\sum_{k}G(s_{j},\tilde{x}_{k})\tilde{w}_{k}-\tilde{u}_{j}\right|^{2}.$
@@
For an actual problem, the number of spikes $n_{x}$ is not known a priori. An important question is to pick the right degree $d$ of the polynomial $p(x)$. The general criteria are that the objective value of (3) (after postprocessing) should be within the noise level, and the degree $d$ should be as small as possible. Commonly used criteria include AIC [akaike1998information] and BIC [schmidt1986multiple].
@@ 2.1. Main idea
Prony’s method uses two special features of the line-spectrum estimation problem: (a) the kernel is of the exponential form, and (b) $\{s_{j}\}$ forms an equally spaced grid. These two features together allow one to write down $M$ (the shift matrix) explicitly. However, these two features no longer hold for sparse recovery problems with a general kernel $G(\cdot,\cdot)$ or unstructured sample locations $\{s_{j}\}$.
@@
Below, we detail how to apply the eigenmatrix idea to complex and real cases. Since the postprocessing step and the choice of the degree $d$ are common to Prony’s method, we present the algorithm for a fixed $d$ in order to simplify the presentation.
@@ 2.2. Complex analytic case
To simplify the discussion, assume first that $X$ is the unit disc $\mathbb{D}_{1}$, and we will comment on the general case at the end. Define for each $x$ the vector $\mathbf{g}(x):=[G(s_{j},x)]_{1\leq j\leq n_{s}}$. The first step is to construct $M$ such that $M\mathbf{g}(x)\approx x\mathbf{g}(x)$ for $x\in\mathbb{D}_{1}$.
Numerically, it is more robust to use the normalized vector $\hat{\mathbf{g}}(x)=\mathbf{g}(x)/\|\mathbf{g}(x)\|$ since the norm of $\mathbf{g}(x)$ can vary significantly depending on $x$. The condition then becomes
$M\hat{\mathbf{g}}(x)\approx x\hat{\mathbf{g}}(x),\quad x\in\mathbb{D}_{1}.$
We enforce this condition on a dense uniform grid $\{a_{k}\}_{1\leq k\leq n_{a}}$ of size $n_{a}$ on the boundary of the unit disk
$M\hat{\mathbf{g}}(a_{k})\approx a_{k}\hat{\mathbf{g}}(a_{k}).$
Define an $n_{s}\times n_{a}$ matrix $W=[\hat{\mathbf{g}}(a_{k})]_{1\leq k\leq n_{a}}$ with $\hat{\mathbf{g}}(a_{k})$ as columns and also an $n_{a}\times n_{a}$ diagonal matrix $\Lambda=\mathrm{diag}(a_{k})$. The above condition can be written in a matrix form as
$MW\approx W\Lambda.$
This suggests the following choice of the eigenmatrix
(4) $M:=W\Lambda W^{\dagger},$
where the pseudoinverse $W^{\dagger}$ is computed by thresholding the singular values of $W$ below $\sigma$.
With the eigenmatrix $M$ available, the rest of the algorithm is similar. Define the vector $\tilde{\mathbf{u}}:=[\tilde{u}_{j}]_{1\leq j\leq n_{s}}$ from the noisy sample values. Since $(M-x_{k}I)\mathbf{g}(x_{k})\approx 0$,
$\prod_{\ell}(M-x_{\ell}I)\mathbf{g}(x_{k})\approx 0\quad\text{and}\quad\prod_{\ell}(M-x_{\ell}I)\tilde{\mathbf{u}}\approx 0.$
Therefore, we can set up the following system:
$\begin{bmatrix}\tilde{\mathbf{u}}&M\tilde{\mathbf{u}}&\ldots&M^{d}\tilde{\mathbf{u}}\end{bmatrix}\tilde{\mathbf{p}}=0\quad\text{with}\quad\tilde{\mathbf{p}}=\begin{bmatrix}\tilde{p}_{0}\\ \vdots\\ \tilde{p}_{d}\end{bmatrix},$
and solve for $\tilde{\mathbf{p}}$ as the null space vector. The roots of the polynomial $\tilde{p}(x):=\tilde{p}_{0}+\tilde{p}_{1}x+\cdots+\tilde{p}_{d}x^{d}$ give the estimators $\{\tilde{x}_{k}\}$ for $\{x_{k}\}$. With $\{\tilde{x}_{k}\}$ available, the least square solve
gives the estimators $\{\tilde{w}_{k}\}$ for $\{w_{k}\}$.
@@
The first question to be answered is why the uniform grid $\{a_{k}\}$ on the unit circle is enough. The following calculation shows why. $M\mathbf{g}(a_{k})\approx a_{k}\mathbf{g}(a_{k})$ implies $M\mathbf{g}(x)\approx x\mathbf{g}(x)$ for all $x\in\mathbb{D}_{1}$:
@@
where the first and third approximations use the exponential convergence of the trapezoidal rule for analytic functions $\mathbf{g}(x)$ and $x\mathbf{g}(x)$, and the second approximation directly comes from $M\mathbf{g}(a_{k})\approx a_{k}\mathbf{g}(a_{k})$. The equalities are applications of the Cauchy integral theorem.
@@
The second question is, what if $X$ is not $\mathbb{D}_{1}$? For a general connected domain $X$ with smooth boundary, let $\phi(t):\mathbb{D}_{1}\rightarrow X$ be the one-to-one map between $\mathbb{D}_{1}$ and $X$ from the Riemann mapping theorem. We then consider the new kernel $G(s,t)=G(s,\phi(t))$ between $S$ and $\mathbb{D}_{1}$ and use the above algorithm to recover the locations $\{\tilde{t}_{k}\}$ in $\mathbb{D}_{1}$. Once $\{\tilde{t}_{k}\}$ are available, we set $\tilde{x}_{k}=\phi(\tilde{t}_{k})$.

### 08 Mar 2024 v4

@@ Abstract.
This note considers the unstructured sparse recovery problems in a general form. Examples include rational approximation, spectral function estimation, Fourier inversion, Laplace inversion, and sparse deconvolution. The main challenges are the noise in the sample values and the unstructured nature of the sample locations. This note proposes the eigenmatrix, a data-driven construction with desired approximate eigenvalues and eigenvectors. The eigenmatrix offers a new way for these sparse recovery problems. Numerical results are provided to demonstrate the efficiency of the proposed method.
@@
Sparse recovery, Prony’s method, ESPRIT algorithm.
@@
This work is partially supported by NSF grants DMS-2011699 and DMS-2208163. The author thanks Laurent Demanet and Lin Lin for helpful discussions and the anonymous referee for constructive comments.
@@ 1. Introduction
This note considers the unstructured sparse recovery problems of a general form. Let $X$ be the parameter space, typically a subset of $\mathbb{R}$ or $\mathbb{C}$, and $S$ be the sampling space. $G(s,x)$ is the kernel function for $s\in S$ and $x\in X$, and is assumed to be analytic in $x$. Suppose that
@@
  • Laplace inversion. $G(s,x)=x\exp(-sx)$, $X$ is an interval $[c_{1},c_{2}]$ of the positive real axis, and $\{s_{j}\}$ is a set of positive real numbers.
@@
The primary challenges of the current setup come from two sources. First, the kernel $G(s,x)$ can be quite general. Second, the sample locations $\{s_{j}\}$ are unstructured, which excludes many existing algorithms that exploit special structures. Third, the sample values $\{\tilde{u}_{j}\}$ are noisy, which raises stability issues when the recovery problem is quite ill-posed.
@@ 1.1. Contribution
This note introduces the eigenmatrix for these unstructured sparse recovery problems. By defining the vector-valued function $\mathbf{g}(x)=[G(s_{j},x)]_{1\leq j\leq n_{s}}$ for $x\in X$, we introduce the eigenmatrix as an $n_{s}\times n_{s}$ matrix $M$ that satisfies, for any $x\in X$
@@ 1.2. Related work
The rest of the note is organized as follows. Section 2 reviews Prony’s method and the ESPRIT algorithms for the special case of the exponential kernel with the uniform sampling grid. Section 3 describes the eigenmatrix approach for the general kernels and unstructured grids. Section 4 presents the numerical experiments of the applications mentioned above. Section 5 concludes with a discussion for future work.
2. Prony and ESPRIT
@@
To motivate the eigenmatrix construction, we first briefly review Prony’s method and the ESPRIT algorithm. Consider the recovery problem with $X=\mathbb{S}\equiv\{z:|z|=1\}\subset\mathbb{C}$, $S=\mathbb{R}$, $G(s,x)=x^{s}$ (with the branch cut at $x=-1$), and $s_{j}=j$ for $j\in\mathbb{Z}$. Here, we make the simplifying assumption that $\{s_{j}\}$ is the whole integer lattice, though only a finite chunk is required in the actual implementations. Most presentations of Prony’s method and the ESPRIT algorithm start with the Hankel matrix. However, our presentation emphasizes the role of the shifting operator in order to motivate the eigenmatrix approach in Section 3.
@@
Introduce the infinitely-long vector-valued function $\mathbf{g}(x)=[G(s_{j},x)]_{j\in\mathbb{Z}}=[x^{j}]_{j\in\mathbb{Z}}$. Let $M$ be the shifting operator that moves each entry up by one slot, i.e., $(M\mathbf{v})_{j}=\mathbf{v}_{j+1}$ for any vector $\mathbf{v}$. Then
$M\mathbf{g}(x)=x\mathbf{g}(x).$
@@
Define the vector $\mathbf{u}=[u_{j}]_{j\in\mathbb{Z}}$ of the exact observations $u_{j}=u(s_{j})=\sum_{k}G(s_{j},x_{k})w_{k}=\sum_{k}w_{k}x_{k}^{j}$. Since $\mathbf{u}=\sum_{k}\mathbf{g}(x_{k})w_{k}$, we have for any $t\geq 0$
$M^{t}\mathbf{u}=\sum_{k}\mathbf{g}(x_{k})w_{k}x_{k}^{t}.$
For the Prony’s method, consider the matrix
$\begin{bmatrix}\mathbf{u}&M\mathbf{u}&\ldots&M^{n_{x}}\mathbf{u}\end{bmatrix}=\begin{bmatrix}\mathbf{g}(x_{1})&\ldots&\mathbf{g}(x_{n_{x}})\end{bmatrix}\begin{bmatrix}w_{1}&&\\ &\ddots&\\ &&w_{n_{x}}\end{bmatrix}\begin{bmatrix}1&x_{1}&\ldots&x_{1}^{n_{x}}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n_{x}}&\ldots&x_{n_{x}}^{n_{x}}\end{bmatrix}$
Let $\mathbf{p}$ be a non-zero vector in its null space, i.e.,
@@
(1) $\begin{bmatrix}\mathbf{u}&M\mathbf{u}&\ldots&M^{n_{x}}\mathbf{u}\end{bmatrix}\mathbf{p}=0\quad\text{with}\quad\mathbf{p}=\begin{bmatrix}p_{0}\\ \vdots\\ p_{n_{x}}\end{bmatrix}.$
Therefore,
$\begin{bmatrix}1&x_{1}&\ldots&x_{1}^{n_{x}}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n_{x}}&\ldots&x_{n_{x}}^{n_{x}}\end{bmatrix}\begin{bmatrix}p_{0}\\ \vdots\\ p_{n_{x}}\end{bmatrix}=0.$
This implies that $\{x_{k}\}$ are the roots of $p(x)=p_{0}+p_{1}x+\ldots p_{n_{x}}x^{n_{x}}$. Therefore, one can identify $\{x_{k}\}$ via rootfinding once $\mathbf{p}$ is computed.
For the ESPRIT algorithm, consider the matrix
$\begin{bmatrix}\mathbf{u}&M\mathbf{u}&\ldots&M^{\ell}\mathbf{u}\end{bmatrix}=\begin{bmatrix}\mathbf{g}(x_{1})&\ldots&\mathbf{g}(x_{n_{x}})\end{bmatrix}\begin{bmatrix}w_{1}&&\\ &\ddots&\\ &&w_{n_{x}}\end{bmatrix}\begin{bmatrix}1&x_{1}&\ldots&x_{1}^{\ell}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n_{x}}&\ldots&x_{n_{x}}^{\ell}\end{bmatrix}$
with $\ell>n_{x}$. Let $USV^{*}$ be the rank-$n_{x}$ singular value decomposition (SVD) of this matrix. The matrix $V^{*}$ takes the form
$V^{*}=P\begin{bmatrix}1&x_{1}&\ldots&x_{1}^{\ell}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n_{x}}&\ldots&x_{n_{x}}^{\ell}\end{bmatrix}$
where $P$ is an unknown non-degenerate $n_{x}\times n_{x}$ matrix. Let $Z_{0}$ and $Z_{1}$ be the submatrices of $V^{*}$ by excluding the first column and the last column, respectively, i.e.,
$Z_{0}=P\begin{bmatrix}1&\ldots&x_{1}^{\ell-1}\\ \vdots&\ddots&\vdots\\ 1&\ldots&x_{n_{x}}^{\ell-1}\end{bmatrix},\quad Z_{1}=P\begin{bmatrix}x_{1}&\ldots&x_{1}^{\ell}\\ \vdots&\ddots&\vdots\\ x_{n_{x}}&\ldots&x_{n_{x}}^{\ell}\end{bmatrix}.$
By introducing
$Z_{1}(Z_{0})^{+}=P\begin{bmatrix}x_{1}&&\\ &\ddots&\\ &&x_{n_{x}}\end{bmatrix}P^{-1},$
one can identify $\{x_{k}\}$ by computing the eigenvalues of $Z_{1}(Z_{0})^{+}$.
@@
For both methods, given $\{x_{k}\}$, the sample weights $\{w_{k}\}$ can be computed via, for example, the least square solve.
@@
For most problems, the sample values $\{\tilde{u}_{j}\}$ have noise. As a result, the sample locations and weights obtained above are only approximations. Many implementations of the Prony and ESPRIT methods have a postprocessing step, where these approximations are used as the initial guesses of the following optimization problem
(2) $\min_{\tilde{x}_{k},\tilde{w}_{k}}\sum_{j}\left|\sum_{k}\tilde{x}_{k}^{j}\tilde{w}_{k}-\tilde{u}_{j}\right|^{2}.$
@@
For an actual problem, the number of spikes $n_{x}$ is not known a priori. An important question is how to pick the right degree of the polynomial $p(x)$ (for Prony’s method) or the rank of the truncated SVD (for the ESPRIT algorithm). The general criteria are that the objective value of (2) (after postprocessing) should be within the noise level, and the degree $d$ should be as small as possible. Commonly used criteria include AIC [akaike1998information] and BIC [schmidt1986multiple].
3. Eigenmatrix
@@ 3.1. Main idea
The discussion above uses two special features of the problem: (a) the kernel is of the exponential form, and (b) $\{s_{j}\}$ forms an equally spaced grid. These two features together allow one to write down the shifting operator $M$ explicitly. However, these two features no longer hold for sparse recovery problems with a general kernel $G(\cdot,\cdot)$ or unstructured sample locations $\{s_{j}\}$.
@@
Below, we detail how to apply the eigenmatrix idea to complex and real cases. Since the choice of degree (in Prony’s algorithm) and the rank of truncated SVD (in the ESPRIT algorithm) remain unchanged, we present the algorithm for fixed $n_{x}$ in order to simplify the discussion.
@@ 3.2. Complex analytic case
To simplify the discussion, assume first that $X$ is the unit disc $\mathbb{D}$, and we will comment on the general case at the end. Define for each $x$ the vector $\mathbf{g}(x):=[G(s_{j},x)]_{1\leq j\leq n_{s}}$. The first step is to construct $M$ such that $M\mathbf{g}(x)\approx x\mathbf{g}(x)$ for $x\in\mathbb{D}$. Numerically, it is more robust to use the normalized vector $\hat{\mathbf{g}}(x)=\mathbf{g}(x)/\|\mathbf{g}(x)\|$ since the norm of $\mathbf{g}(x)$ can vary significantly depending on $x$. The condition then becomes
$M\hat{\mathbf{g}}(x)\approx x\hat{\mathbf{g}}(x),\quad x\in\mathbb{D}.$
We enforce this condition on a uniform grid $\{a_{t}\}_{1\leq t\leq n_{a}}$ of size $n_{a}$ on the boundary of the unit disk
$M\hat{\mathbf{g}}(a_{t})\approx a_{t}\hat{\mathbf{g}}(a_{t}).$
Define an $n_{s}\times n_{a}$ matrix $\hat{G}=[\hat{\mathbf{g}}(a_{t})]_{1\leq t\leq n_{a}}$ with $\hat{\mathbf{g}}(a_{t})$ as columns and also an $n_{a}\times n_{a}$ diagonal matrix $\Lambda=\mathrm{diag}(a_{t})$. The above condition can be written in a matrix form as
(3) $M\hat{G}\approx\hat{G}\Lambda.$
@@
The main guideline for the choice of $n_{a}$ is that the columns of $\hat{G}$ are numerically linearly independent. To see why this is essential, let us consider the extreme case of a kernel $G(s,x)$ constant in $x$. Here, the columns are linearly dependent, and there is no way to recover individual $\{x_{k}\}$. In practice, different sampling locations $\{s_{j}\}$ leads to different choices of $n_{a}$. In practice, $n_{a}$ is chosen such that the condition number of $\hat{G}$ is bounded below $10^{7}$.
When the columns of $\hat{G}$ are numerically linearly independent, (3) suggests the following choice of the eigenmatrix
(4) $M:=\hat{G}\Lambda\hat{G}^{\dagger},$
where the pseudoinverse $\hat{G}^{\dagger}$ is computed by thresholding the singular values of $\hat{G}$. Since $M$ is to be applied repetitively as in (1), the threshold is set so that the norm of $M$ is bounded by a small constant such as 3.
Remark 4.
A key question is why enforcing the condition on the uniform grid $\{a_{t}\}$ is enough. The following calculation shows why. $M\mathbf{g}(a_{t})\approx a_{t}\mathbf{g}(a_{t})$ implies $M\mathbf{g}(x)\approx x\mathbf{g}(x)$ for all $x\in\mathbb{D}$:
@@
where the first and third approximations use the exponential convergence of the trapezoidal rule for analytic functions $\mathbf{g}(x)$ and $x\mathbf{g}(x)$, and the second approximation directly comes from $M\mathbf{g}(a_{t})\approx a_{t}\mathbf{g}(a_{t})$. The equalities are applications of the Cauchy integral theorem.
@@
The second question is, what if $X$ is not $\mathbb{D}$? For a general connected domain $X$ with smooth boundary, let $\phi(t):\mathbb{D}\rightarrow X$ be the one-to-one map between $\mathbb{D}$ and $X$ from the Riemann mapping theorem. We then consider the new kernel $G(s,t)=G(s,\phi(t))$ between $S$ and $\mathbb{D}$ and use the above algorithm to recover the locations $\{\tilde{t}_{k}\}$ in $\mathbb{D}$. Once $\{\tilde{t}_{k}\}$ are available, we set $\tilde{x}_{k}=\phi(\tilde{t}_{k})$.
To provide an idea about the eigenmatrix, we apply the above construction to the problem given in Section 2 with $n_{s}=32$. Figure 1 plots the eigenmatrix for three cases: (a) $\{s_{j}=j\}$ is a integer lattice (the classical case); (b) $\{s_{j}\}$ is a random perturbation of the integer lattice; (c) $\{s_{j}\}$ are chosen uniformly in $[0,n_{s}]$. In the integer lattice case, the eigenmatrix approach reproduces the shifting matrix. In the perturbative case, the eigenmatrix is quite close to the shifting matrix, as expected. In the fully random case, the eigenmatrix is quite different but still keeps the overall trend.
Figure 1. The eigenmatrix for the example in Section 2. $X=\mathbb{S}\equiv\{z:|z|=1\}\subset\mathbb{C}$, $S=\mathbb{R}$, $G(s,x)=x^{s}$ (with the branch cut at $x=-1$). $n_{s}=32$. Left: $M$ when $s_{j}=j$ for $0\leq j<n_{s}$. Middle: $M$ when $s_{j}$ is a random perturbation of the integer lattice. Right: $M$ when $\{s_{j}\}$ are chosen uniformly in $[0,n_{s}]$.